# K3 surfaces with real or complex multiplication

### v2

@@ 12.1. Proof of Theorem B
Given a CM field $E$ of degree $d$ over $\mathbf{C}$ and an integer $m$ such that $md\leq 20$, Proposition 6.4 provides a quadratic form $W$ over $E$ such that ${\mathrm{T}}(W)$ has signature $(2,md-2)$ and embeds into $V_{K3}$. Applying Theorem 11.2 to $\mathrm{T}(W)$ with its natural $E$-module structure, we obtain an $(m-1)$-dimensional family of complex projective K3 surfaces with $E={\rm End}_{\rm Hdg}(T_{X,{\mathbf{Q}}})$ for a very general member $X$.
@@ 12.2. Proof of Theorem A
Let $E$ be totally real of degree $d$ and $m\geq 3$. From Corollary 7.4 we obtain $W$ over $E$ such that $V_{K3}=\mathrm{T}(W)\oplus V^{\prime}$ and the $dm$-dimensional quadratic form $\mathrm{T}(W)$, which satisfies condition (C) by construction, also satisfies the other conditions in Theorem 11.2. Hence there is an $(m-2)$-dimensional family of complex projective K3 surfaces with $E={\rm End}_{\rm Hdg}(T_{X,{\mathbf{Q}}})$ for a very general member $X$. ∎
@@
Theorem A (and B) ‘only’ show that for any totally real (or CM) field $E$ there exists a K3 surface $\tilde{X}$, with a transcendental lattice of a given rank, such that $E={\rm End}_{\rm Hdg}(T_{\tilde{X}})$. However, Proposition 7.3 actually gives a much more precise result, showing that the Picard lattice $\mathrm{Pic}(\tilde{X})$, of rank one, can be chosen arbitrarily.
@@ 13. Picard lattices
Suppose that $E$ is totally real and $m\geqslant 3$. Assume moreover that $d$ is odd or ${\rm det}(L)\Delta_{E}\in\Lambda^{+}_{E/{{\mathbf{Q}}}}$. Then there exists a family of complex projective $K3$ surfaces $X$ such that a very general member of this family satisfies
$\bullet$ $A_{X}\simeq E$;
$\bullet$ ${\rm Pic}(X)\simeq L$.
Proof. Note that $V^{\prime}:=L\otimes_{\bf Z}{{\mathbf{Q}}}$ embeds into $V_{K3}=\Lambda_{K3}\otimes_{\bf Z}{{\mathbf{Q}}}$. Set $V=V_{K3}$, and let $U$ be the orthogonal complement of $V^{\prime}$ in $V$, i.e. $U\oplus V^{\prime}\simeq V$. By Theorem 9.3 (if $d$ is odd) and Theorem 9.5 (if $d$ is even) there exists a quadratic form $W$ over $E$ such that $U=T(W)$; moreover, there is an embedding $\sigma\in\Sigma_{E}$ such that the signature of $W_{\sigma}$ is $(2,m-2)$. The result for the quadratic form $V^{\prime}=L\otimes{\mathbf{Q}}$ thus follows from Theorem 11.2. The statement for the precise given lattice structure then follows from the given primitive embedding $L\hookrightarrow\Lambda_{K3}$, as pointed out in Section 10.6. ∎
@@
Suppose that $E$ is a CM field and that $m\geqslant 2$. There exists a family of complex projective $K3$ surfaces $X$ such that a very general member of this family satisfies
$\bullet$ $A_{X}\simeq E$;
$\bullet$ ${\rm Pic}(X)\simeq L$
if and only if ${\rm disc}(V^{\prime})=\Delta_{E}$, and $L\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p}$ is isomorphic to an orthogonal sum of hyperbolic planes for all $p\in S_{E}$.
Proof. As in the previous proof, let $V=V_{K3}$, and let $U$ be the orthogonal complement of $V^{\prime}=L\otimes_{\bf Z}{{\mathbf{Q}}}$ in $V$, i.e. $U\oplus V^{\prime}\simeq V$. Since ${\mathrm{disc}}(V)=1$, we have ${\mathrm{disc}}(U)=\Delta_{E}$. If $p\in S_{E}$, then $V^{\prime}\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p}$ is isomorphic to a sum of hyperbolic planes. This implies that $U\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p}$ is also isomorphic to a sum of hyperbolic planes. The signature of $U$ is $(2,md-2)$. Applying Theorem 9.2 we conclude that there exists a hermitian form $W$ such that $U\simeq T(W)$. Then we can use Theorem 11.2 to conclude the existence of the desired K3 surface. ∎
@@ 13.1. Connection with elliptic fibrations
Proof. Let $E^{\prime}\subset A_{X}$ be a CM field of degree $d=2$ or $10$, then $m=10$ or $2$ and hence $\Delta_{E}^{m}$ is a square. Then by Proposition 6.4 we have $\mathrm{Pic}(X)\otimes{\mathbf{Q}}=\mathrm{T}(W)^{\perp}=H$. In particular, $\mathrm{Pic}(X)$ represents zero. It is well known that this implies that $X$ admits an elliptic fibration as stated (i.e. almost all fibres are elliptic curves, but note that the fibration may not admit a section, cf. [SS 19, Thm. 11.24]). ∎
@@
For $E=A_{X}$ imaginary quadratic, an explicit construction of an $E$-module structure on $T_{X,{\mathbf{Q}}}$ was given in [vGS 23, Thm. 3.10]. (That construction is also valid for real quadratic fields.)
@@
In contrast, if one considers a K3 surface $X$ admitting a non-symplectic automorphism of order $5$, this gives CM by ${\mathbf{Q}}(\zeta_{5})$ which is not covered by Proposition 13.4. Indeed, then the maximal family from [AST 11] has dimension $4$, which corresponds to $m=5$ in Theorem 11.2. A very general member has Picard lattice of rank two and determinant $-5$, thus not representing zero and in particular not allowing for a genus one fibration.
@@
Conversely, if $\Delta_{E}$ is not a square, then neither is $-\det(\mathrm{Pic}(X))$; in particular, $\mathrm{Pic}(X)$ cannot represent zero, hence $X$ does not have any elliptic fibration. ∎
@@ 14.1. HK with RM
Then there exists an $m-2$-dimensional family of projective HK manifolds with $b_{2}=r$, of any dimension $n\geq 1$ in case $r=7,23$, such that the very general member $X$ has the properties
$A_{X}\simeq E,\quad\mathrm{dim}_{E}(T_{X,{\mathbf{Q}}})=m,\quad\mathrm{rk}(\mathrm{Pic}(X))=r-md\leavevmode\nobreak\ .$
Proof. In case $r=7$, a $2n$-dimensional HK manifold of generalized Kummer type has $H^{2}(X,{\mathbf{Q}})\simeq H^{3}\oplus\langle-2n-2\rangle$. Proposition 7.5 shows that if there is an isomorphism of lattices $\mathrm{Pic}(X)\simeq\langle h\rangle$ for some $h\in 2\mathbf{N}$ with $-2(n+1)h\in{\mathrm{N}}_{E/{\mathbf{Q}}}(E^{\times})$, then $T_{X}=\mathrm{T}(W)$ for a quadratic form $W$ over $E$ with the required signature Condition (C) from Definition 11.1. As $\dim T_{X,{\mathbf{Q}}}=6=2\cdot 3$, Theorem 11.2 gives an $3-2=1$ dimensional family of such manifolds with RM by a quadratic field $E$, the very general member of which has the desired properties.
@@
In case $r=23$ and $\mathrm{Pic}(X)=\langle h\rangle$, $T_{X,{\mathbf{Q}}}$ has dimension $22$ and thus if $A_{X}$ is a totally real field $E$, it must have degree $2$ and $\dim_{E}(T_{X,{\mathbf{Q}}})=11$. Proposition 7.5 shows that if $-2(n-1)h\in{\mathrm{N}}_{E/{\mathbf{Q}}}(E^{\times})$ then $T_{X}=\mathrm{T}(W)$ for a quadratic form $W$ over $E$ with the required signature condition and we obtain an $11-2=9$-dimensional family of such manifolds with RM by the quadratic field $E$, again from Theorem 11.2. In case $\mathrm{rk}(\mathrm{Pic}(X))>1$, the result follows again from the combination of Corollary 7.2 and Theorem 11.2.
@@ 14.2. HK with CM
Then there exists an $(m-1)$-dimensional family of projective HK manifolds with $b_{2}=r$, of any dimension $n\geq 1$ in case $r=7,23$, such that the very general member $X$ has the properties
$A_{X}\simeq E,\quad\mathrm{dim}_{E}(T_{X,{\mathbf{Q}}})=m,\quad\mathrm{rk}(\mathrm{Pic}(X))=r-md\leavevmode\nobreak\ .$
@@ 14.3. The Hodge group and the Mumford–Tate group
Zarhin showed that the Hodge group, also known as the Special Mumford-Tate group of a simple rational Hodge structure K3 type $V$ is determined by its endomorphism algebra $E$, recall that there is then a form $W$ over $E$ (quadratic or hermitian) such that there is an isometry $V\simeq\mathrm{T}(W)$. In fact, the Hodge group is the subgroup of $E$-linear special isometries ${\rm SO}_{E}(V)\subset SO(V)$ [Z 83, Thm. 2.2.1, Thm. 2.3.1], cf. [H 16], Chap. 3, §3, especially Theorem 3.9.
The following result shows that if $md<20$, every such special orthogonal and unitary group occurs for some K3 surface.
@@
Let $E$ be a totally real or CM field of degree $d$ and let $m\geqslant 1$ be an integer such that $md<20$. Let $W$ be a quadratic (respectively hermitian) form of dimension $d$ over $E$ such that $\mathrm{T}(W)$ has signature $(2,md-2)$. If $W$ is quadratic, assume in addition that $m\geqslant 3$ and that there exists a real embedding of $E$ such that $W_{\sigma}$ has signature $(2,m-2)$. Then there exists a complex projective $K3$ surface with Hodge group isomorphic to $\Pi_{E/{{\mathbf{Q}}}}({\rm SO}(W))$ if $E$ is totally real and $\Pi_{E/{{\mathbf{Q}}}}({\rm U}(W))$ if $E$ is a CM field.
@@ Acknowledgements
We thank Simon Brandhorst and John Voight for very helpful comments.

### v3

@@ 13.1. Proof of Theorem B
Given a CM field $E$ of degree $d$ over $\mathbf{C}$ and an integer $m$ such that $md\leq 20$, Proposition 7.4 provides a quadratic form $W$ over $E$ such that ${\mathrm{T}}(W)$ has signature (2,md2)(2,md-2) and embeds into $V_{K3}$. Applying Theorem 12.2 to $\mathrm{T}(W)$ with its natural $E$-module structure, we obtain an (m1)(m-1)-dimensional family of complex projective K3 surfaces with $E={\rm End}_{\rm Hdg}(T_{X,{\mathbf{Q}}})$ for a very general member $X$.
@@ 13.2. Proof of Theorem A
Let $E$ be totally real of degree $d$ and $m\geq 3$. From Corollary 8.4 we obtain $W$ over $E$ such that $V_{K3}=\mathrm{T}(W)\oplus V^{\prime}$ and the $dm$-dimensional quadratic form $\mathrm{T}(W)$, which satisfies condition (C) by construction, also satisfies the other conditions in Theorem 12.2. Hence there is an (m2)(m-2)-dimensional family of complex projective K3 surfaces with $E={\rm End}_{\rm Hdg}(T_{X,{\mathbf{Q}}})$ for a very general member $X$. ∎
@@
[…]
@@ 14. Picard lattices
Suppose that $E$ is totally real and $m\geqslant 3$. Assume moreover that $d$ is odd or ${\rm det}(L)\Delta_{E}\in\Lambda^{+}_{E/{{\mathbf{Q}}}}$. Then there exists an (m2)(m-2)-dimensional family of complex projective $K3$ surfaces $X$ such that a very general member of this family satisfies
  • $\bullet$
    $A_{X}\simeq E$;
  • $\bullet$
    ${\rm Pic}(X)\simeq L$.
Proof. Note that $V^{\prime}:=L\otimes_{\bf Z}{{\mathbf{Q}}}$ embeds into $V_{K3}=\Lambda_{K3}\otimes_{\bf Z}{{\mathbf{Q}}}$. Set $V=V_{K3}$, and let $U$ be the orthogonal complement of $V^{\prime}$ in $V$, i.e. $U\oplus V^{\prime}\simeq V$. By Theorem 10.4 (if $d$ is odd) and Theorem 10.6 (if $d$ is even) there exists a quadratic form $W$ over $E$ such that $U=T(W)$; moreover, there is an embedding $\sigma\in\Sigma_{E}$ such that the signature of $W_{\sigma}$ is (2,m2)(2,m-2). The result for the quadratic form $V^{\prime}=L\otimes{\mathbf{Q}}$ thus follows from Theorem 12.2. The statement for the precise given lattice structure then follows from the given primitive embedding $L\hookrightarrow\Lambda_{K3}$, as pointed out in Section 11.6. ∎
@@
Suppose that $E$ is a CM field and that $m\geqslant 2$. There exists an (m1)(m-1)-dimensional family of complex projective $K3$ surfaces $X$ such that a very general member of this family satisfies
  • $\bullet$
    $A_{X}\simeq E$;
  • $\bullet$
    ${\rm Pic}(X)\simeq L$
if and only if ${\rm disc}(L\otimes_{\bf Z}{{\mathbf{Q}}})=\Delta_{E}^{m}$, and $L\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p}$ is isomorphic to an orthogonal sum of hyperbolic planes for all $p\in S_{E}$.
Proof. The conditions are necessary by Theorem 10.2. Let us prove that they are sufficient. As in the previous proof, let $V=V_{K3}$, and let $U$ be the orthogonal complement of $V^{\prime}=L\otimes_{\bf Z}{{\mathbf{Q}}}$ in $V$, i.e. $U\oplus V^{\prime}\simeq V$. Since ${\mathrm{disc}}(V)=1$, we have ${\mathrm{disc}}(U)=\Delta_{E}^{m}$. If $p\in S_{E}$, then $V^{\prime}\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p}$ is isomorphic to a sum of hyperbolic planes. This implies that $U\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p}$ is also isomorphic to a sum of hyperbolic planes. The signature of $U$ is (2,md2)(2,md-2). Applying Theorem 10.2 we conclude that there exists a hermitian form $W$ such that $U\simeq T(W)$. Then we can use Theorem 12.2 to conclude the existence of the desired family of K3 surfaces. ∎
@@ 14.1. Connection with elliptic fibrations
Proof. Let $E\subset A_{X}$ be a CM field of degree $d=2$ or $10$, then $m=10$ or $2$ and hence $\Delta_{E}^{m}$ is a square. Then by Proposition 7.4 we have $\mathrm{Pic}(X)\otimes{\mathbf{Q}}=\mathrm{T}(W)^{\perp}=H$. In particular, $\mathrm{Pic}(X)$ represents zero. It is well known that this implies that $X$ admits an elliptic fibration as stated (i.e. almost all fibres are elliptic curves, but note that the fibration may not admit a section, cf. [SS 19, Thm. 11.24]). ∎
We also record the following converse of Proposition 14.4 on the level of lattices:
Lemma 14.5.
Let $X$ be an elliptic K3 surface with $\rho(X)=2$ and $E$ a CM field of degree $2$ or $10$. Then $L=\mathrm{NS}(X)$ and $E$ satisfy the conditions of Theorem 14.3.
Proof. Since disc(L)=1(L)=1 and $m$ is even, this amounts to an easy check using first the elementary fact that $L\simeq H$ over $\mathbf{Q}$} (derived from the existence of a non-zero isotropic vector in $L$), so $L^{\perp}\simeq H^{2}\oplus E_{8}^{2}$, and second the classification of $p$-adic quadratic forms to verify that $E_{8}\simeq H^{4}$ over ${\mathbf{Q}}_{p}$ for any prime $p$. ∎
@@
For $E$ imaginary quadratic, this validates the claim from [vGS 25, Rem. 3.13], even though the $E$-module structure on $L^{\perp}$ given there is not sufficient as it is not compatible with the complex conjugation condition (0.1).
@@
In contrast, if one considers a K3 surface $X$ admitting a non-symplectic automorphism of order $5$, this gives CM by ${\mathbf{Q}}(\zeta_{5})$ which is not covered by Proposition 14.4. Indeed, then the maximal family from [AST 11] has dimension $4$, which corresponds to $m=5$ in Theorem 12.2. A very general member has Picard lattice of rank two and determinant 5-5, thus not representing zero and in particular not allowing for a genus one fibration.
@@
Conversely, if $\Delta_{E}$ is not a square, then neither is $\det(\mathrm{Pic}(X))$; in particular, $\mathrm{Pic}(X)$ cannot represent zero, hence $X$ does not have any elliptic fibration. ∎
@@ 15.1. HK with RM
Then there exists an m2m-2-dimensional family of projective HK manifolds with $b_{2}=r$, of any dimension $n\geq 1$ in case $r=7,23$, such that the very general member $X$ has the properties
$A_{X}\simeq E,\quad\mathrm{dim}_{E}(T_{X,{\mathbf{Q}}})=m,\quad\mathrm{rk}(\mathrm{Pic}(X))=r-md\penalty 10000\ .$
Proof. In case $r=7$, a $2n$-dimensional HK manifold of generalized Kummer type has $H^{2}(X,{\mathbf{Q}})\simeq H^{3}\oplus\langle-2n-2\rangle$. Proposition 8.5 shows that if there is an isomorphism of lattices $\mathrm{Pic}(X)\simeq\langle h\rangle$ for some $h\in 2\mathbf{N}$ with $2(n+1)h\in{\mathrm{N}}_{E/{\mathbf{Q}}}(E^{\times})$, then $T_{X}=\mathrm{T}(W)$ for a quadratic form $W$ over $E$ with the required signature Condition (C) from Definition 12.1. As $\dim T_{X,{\mathbf{Q}}}=6=2\cdot 3$, Theorem 12.2 gives an 32=13-2=1 dimensional family of such manifolds with RM by a quadratic field $E$, the very general member of which has the desired properties.
@@
In case $r=23$ and $\mathrm{Pic}(X)=\langle h\rangle$, $T_{X,{\mathbf{Q}}}$ has dimension $22$ and thus if $A_{X}$ is a totally real field $E$, it must have degree $2$ and $\dim_{E}(T_{X,{\mathbf{Q}}})=11$. Proposition 8.5 shows that if $2(n-1)h\in{\mathrm{N}}_{E/{\mathbf{Q}}}(E^{\times})$ then $T_{X}=\mathrm{T}(W)$ for a quadratic form $W$ over $E$ with the required signature condition and we obtain an 112=911-2=9-dimensional family of such manifolds with RM by the quadratic field $E$, again from Theorem 12.2. In case $\mathrm{rk}(\mathrm{Pic}(X))>1$, the result follows again from the combination of Corollary 8.2 and Theorem 12.2.
@@ 15.2. HK with CM
Then there exists an (m1)(m-1)-dimensional family of projective HK manifolds with $b_{2}=r$, of any dimension $n\geq 1$ in case $r=7,23$, such that the very general member $X$ has the properties
$A_{X}\simeq E,\quad\mathrm{dim}_{E}(T_{X,{\mathbf{Q}}})=m,\quad\mathrm{rk}(\mathrm{Pic}(X))=r-md\penalty 10000\ .$
@@ 15.3. The Hodge group and the Mumford–Tate group
Zarhin showed that the Hodge group, also known as the Special Mumford-Tate group, of a simple rational Hodge structure K3 type $V$ is determined by its endomorphism algebra $E$, recall that there is then a form $W$ over $E$ (quadratic or hermitian) such that there is an isometry $V\simeq\mathrm{T}(W)$. In fact, the Hodge group is the subgroup of $E$-linear special isometries ${\rm SO}_{E}(V)\subset SO(V)$ [Z 83, Thm. 2.2.1, Thm. 2.3.1], cf. [H 16], Chap. 3, §3, especially Theorem 3.9.
It is natural to ask which groups actually occur. The following result shows that if $md<20$, every such special orthogonal and unitary group occurs for some K3 surface. To introduce the necessary notation, note that the group $\rm SO$} is an algebraic group defined over the field $E$, and SU is defined over $E_{0}$. To obtain the corresponding Mumford-Tate group, we need to consider these groups as defined over $\mathbf{Q}$}; this is achieved by taking the restriction of scalars, denoted by ResE/Q.
@@
Let $E$ be a totally real or CM field of degree $d$ and let $m\geqslant 1$ be an integer such that $md<20$. Let $W$ be a quadratic (respectively hermitian) form of dimension $d$ over $E$ such that $\mathrm{T}(W)$ has signature (2,md2)(2,md-2). If $W$ is quadratic, assume in addition that $m\geqslant 3$ and that there exists a real embedding of $E$ such that $W_{\sigma}$ has signature (2,m2)(2,m-2). Then there exists a complex projective $K3$ surface with Hodge group isomorphic to ${\rm Res}_{E/{{\mathbf{Q}}}}({\rm SO}(W))$ if $E$ is totally real and ${\rm Res}_{E/{{\mathbf{Q}}}}({\rm U}(W))$ if $E$ is a CM field.
@@ Acknowledgements
We thank Simon Brandhorst and John Voight for very helpful comments. We are very grateful to the referee for many useful comments and corrections.